# Hamiltonian simulation-based quantum-selected configuration interaction for large-scale electronic structure calculations with a quantum computer

### v2

@@ V Conclusion
In this study, we proposed an HSB-QSCI, which uses a quantum computer to perform Hamiltonian simulation and subsequent measurement of the quantum state in the computational basis to sample the important Slater determinants used to span the subspace Hamiltonian to be diagonalized on a classical computer. Compared to the reported VQE-based state preparation approaches for the QSCI [14, 16, 17], the HSB-QSCI is free from variational optimizations of the quantum states, and it works with the simple initial wave function like the HF. Since long-time evolution is in general difficult to simulate on a classical computer, we expect this approach to become powerful as the system size increases. Proof-of-concept simulations were performed for the S0 and T1 states of organic molecules 1–5, and we achieved chemical precision in all molecules studied. The quantum hardware demonstrations of the HSB-QSCI for the S0 and T1 states of three carbyne molecules 5–7 were also reported using ibm_kawasaki with 20, 28, and 36 qubits, respectively, with the aid of MPO-based classical optimization of the quantum circuit for the time evolution operator. The HSB-QSCI achieved chemical precision for the total energy in 5, for the singlet–triplet energy difference in 6, and for the energy difference between planar and twisted geometries of the S0 state in 7. In 7, by considering about 1% of the Slater determinants, the HSB-QSCI captured more than 99.75% of correlation energies in the active space. The largest system studied in this work is 7 with 36 qubits, which is the largest system capable of performing the CAS-CI calculation in a typical computational environment without a supercomputer [45, 55]. We expect that the HSB-QSCI can handle the system larger than 36 qubits, by introducing batch-based implementations in the Hamiltonian matrix diagonalization part, and calculating the variance of the energy expectation value as explored in Reference [19]. Note that, as discussed by Reinholdt and coworkers in their recent paper [56], the QSCI wave function is in some cases less compact than that obtained by classical heuristics. From Eq. (6), the probability that a particular Slater determinant is measured in the HSB-QSCI depends not only on the CI coefficient of the ground state, but also on the gap to the excited state and the CI coefficient of the excited state, and screening out unimportant Slater determinants from the subspace is very important in the application to larger systems. Applying the auxiliary-field quantum Monte Carlo (AFQMC) method with the QSCI wave function as the input can be another direction towards larger-scale quantum chemical calculations with a quantum computer, as suggested in recent studies [57, 58]. Application of the HSB-QSCI to larger systems is currently in progress.

### v3

@@ V Conclusion
In this study, we proposed an HSB-QSCI, which uses a quantum computer to perform Hamiltonian simulation and subsequent measurement of the quantum state in the computational basis to sample the important Slater determinants used to span the subspace Hamiltonian to be diagonalized on a classical computer. Compared to the reported VQE-based state preparation approaches for the QSCI [14, 24, 25], the HSB-QSCI is free from variational optimizations of the quantum states, and it works with the simple initial wave function like the HF. Since long-time evolution is in general difficult to simulate on a classical computer, we expect this approach to become powerful as the system size increases. The fact that higher-order excitations can be incorporated even in a short time evolution also makes HSB-QSCI suitable for simulation of large and strongly correlated systems. Proof-of-concept simulations were performed for the S0 and T1 states of organic molecules 1–5, and we achieved chemical precision in all molecules studied. The quantum hardware demonstrations of the HSB-QSCI for the S0 and T1 states of three carbyne molecules 5–7 were also reported using ibm_kawasaki with 20, 28, and 36 qubits, respectively, with the aid of MPO-based classical optimization of the quantum circuit for the time evolution operator. The HSB-QSCI achieved chemical precision for the total energy in 5, for the singlet–triplet energy difference in 6, and for the energy difference between planar and twisted geometries of the S0 state in 7. In 7, by considering about 1% of the Slater determinants, the HSB-QSCI captured more than 99.18% of correlation energies in the active space. The largest system studied in this work is 7 with 36 qubits, which is the largest system capable of performing the CAS-CI calculation in a typical computational environment without a supercomputer [57, 70]. We expect that the HSB-QSCI can handle the system larger than 36 qubits, by introducing batch-based implementations in the Hamiltonian matrix diagonalization part, and calculating the variance of the energy expectation value as explored in Reference [27]. Note that, as discussed by Reinholdt and coworkers in their recent paper [71], the QSCI wave function is in some cases less compact than that obtained by classical heuristics. From Eq. (7), the probability that a particular Slater determinant is measured in the HSB-QSCI depends not only on the CI coefficient of the ground state, but also on the gap to the excited state and the CI coefficient of the excited state, and screening out unimportant Slater determinants from the subspace is very important in the application to larger systems. Applying the auxiliary-field quantum Monte Carlo (AFQMC) method with the QSCI wave function as the input can be another direction towards larger-scale quantum chemical calculations with a quantum computer, as suggested in recent studies [20, 21]. Application of the HSB-QSCI to larger systems is currently in progress.
@@ VI Acknowledgments
System
HF
CAS-CI
HSB-QSCI
/Hartree
#Dets
/Hartree
%#Detsa
%b
5 (Planar, S0)
63504
100.00
0.00
100.00
5 (Planar, T1)
44100
100.00
0.00
100.00
5 (Twisted, S0)
63504
100.00
0.00
100.00
5 (Twisted, T1)
44100
100.00
0.01
100.00
6 (Planar, S0)
11778624
100.00
0.00
100.00
6 (Planar, T1)
9018009
95.50
0.01
100.00
6 (Twisted, S0)
11778624
99.83
0.00
100.00
6 (Twisted, T1)
9018009
98.57
0.00
100.00
7 (Planar, S0)
2363904400
48.29
0.31
99.99
7 (Planar, T1)
1914762564
27.48
3.70
99.88
7 (Twisted, S0)
2363904400
46.39
0.30
99.99
7 (Twisted, T1)
1914762564
28.96
2.40
99.92
a Percentage of the number of Slater determinants calculated as 100
b Percentage of the correlation energy in the active space considered in the HSB-QSCI, calculated as 100 .